OpenStudy (anonymous):

The volume of a square prism is 84 cubic meters. If the base area is 14meters, What is the length?

OpenStudy (lgbasallote):

i assume by length you mean height?

OpenStudy (anonymous):

Yes sorry

OpenStudy (lgbasallote):

\[V = \text{base area}\times \text{height}\]

OpenStudy (zzr0ck3r):

so we have x^2 = 14 right?

OpenStudy (anonymous):

and that gives me the answer? @lgbasallote

OpenStudy (anonymous):

what? @zzr0ck3r

OpenStudy (zzr0ck3r):

its a square with area 14, how long are the sides?

OpenStudy (lgbasallote):

yes @eAnthonyR

OpenStudy (anonymous):

Answer=1176?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

no...

OpenStudy (lgbasallote):

remember what's given to you are V and base area...you're looking for height

OpenStudy (anonymous):

so 84=14*h?

OpenStudy (anonymous):

So divide both sides right? @lgbasallote

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

6 is the answer

OpenStudy (zzr0ck3r):

im confused

OpenStudy (anonymous):

wait 6 or 6m^2?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (zzr0ck3r):

length will be m

OpenStudy (zzr0ck3r):

If this is a cube, and the base is a square with area 14 how is the length not sqrt(14)?

OpenStudy (anonymous):

Cause its height instead of length man. @zzr0ck3r

OpenStudy (zzr0ck3r):

but its a cube, they should all be the same?

OpenStudy (zzr0ck3r):

o snap I see I read wrong

OpenStudy (lgbasallote):

that was why i asked immediately =)) it seemed fishy

OpenStudy (zzr0ck3r):

I for some reason thought I read cube...sorry

OpenStudy (lgbasallote):

yeah square prism made me think of cube when i first saw them

OpenStudy (anonymous):

the height of a hexagonal prism is 6ft. Its base edges all equal 4ft. Find its volume

OpenStudy (zzr0ck3r):

what is the area of a hexagon?

OpenStudy (zzr0ck3r):

The hexagon has 6 sides and 6 vertices. Label the vertices A, B, C, D, E, and F. Draw a line from A → D, B → E, and C→ F. The lines intersect at the center of the hexagon. You should see 6 triangles. The area of the hexagon = 6 * area of 1 triangle. then multiply that by 6

OpenStudy (zzr0ck3r):

make sense?